OpenStudy (anonymous):

Find the exponential function f(x)=a^x whose graph goes through the point (3, 1/64)?

OpenStudy (bahrom7893):

1/64 = a^3 a = cube root of (1/64) = 1/4

OpenStudy (bahrom7893):

f(x) = (1/4)^x is your function

OpenStudy (anonymous):

Awesome thanks very much!